OpenStudy (anonymous):

Part 1: Explain, in complete sentences, whether or not the expression x2 + 24x + 144 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

OpenStudy (anonymous):

it is a perfect square

OpenStudy (anonymous):

we know because \[144=12^2\] and \[2\times 12=24\]

OpenStudy (anonymous):

that is two point right there. because it is a perfect square it can be factored \[(x+12)^2\] is how you factor.

OpenStudy (anonymous):

i don't understand the distinction between questions 2 and 3, but that is on you